OpenStudy (anonymous):

solve : 14-4y<(with the line ) 66

OpenStudy (johnweldon1993):

\[14 - 4y \le 66\] we want to solve for 'y'...so we isolate it....lets subtract 14 from both sides of the inequality 14 - 4y <= 66 -14 -14 -4y <= 52 Now we want to solve for 'y'...so divide both sides by -4 *Important! when multiplying or dividing by a negative number in an inequality SWITCH THE DIRECTION OF THE SIGN* -4y <= 52 ----------- -4 -4 \[y \ge -13\] Notice the sign went from \[\le\] to \[\ge\]